OpenStudy (anonymous):

Left on together, the cold and hot water faucets of a certain bathtub take 7 minutes to fill the tub. If it takes the hot water faucet 14 minutes to fill the tub by itself, how long will it take the cold water faucet to fill the tub on its own?

OpenStudy (anonymous):

in one minute 1/7 part of bathtub is filled when both cold and hot water left on together( as they fill it in 7 min hot water fill it in 14 min in 1 min --- 1/14 th part of bathtub is filled let time taken my cold water to fill tub is x so in 1 min --- 1/x part of bathtub is filled..

OpenStudy (anonymous):

by above to equation 1/7= 1/14+1/x \[1/ x= 1/7-1/14 \] 1/x=1/7 then total 1/7 part of tank is filled in 1 min so full tub wiil in 7 MINUTE